### v1

# Low complexity binary words avoiding $(5/2)^{+}$-powers

Abstract.
Rote words are infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$. Shallit and Shur, as well as Ollinger and Shallit, showed that there are Rote words that avoid $(5/2)^{+}$-powers and that this is best possible. In this note we give a structure theorem for the Rote words that avoid $(5/2)^{+}$-powers, confirming a conjecture of Ollinger and Shallit.
The work of James Currie is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number DDG-2024-00005].
The work of Narad Rampersad is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2019-04111].
1. Introduction
Two central concepts in combinatorics on words are power avoidance and factor complexity. Recently, Shallit and Shur [9] initiated the systematic investigation of the interplay between these two concepts. They examined two dual problems: 1) Given a particular power avoidance constraint, determine the range of possible factor complexities among all infinite words avoiding the specified power; and, 2) Given a class of words with specified factor complexities, determine the powers that are avoided by at least one word in this class.
A well-known classical result provides a solution for the latter problem for the class of Sturmian words, i.e., the class of infinite words that contain $n+1$ factors of length $n$ for every $n\geq 1$: the Fibonacci word avoids $(5+\sqrt{5})/2$-powers, and this is best possible among all Sturmian words [1]. The Sturmian words are the aperiodic infinite words with the least possible factor complexity function; Shallit and Shur [9] and Ollinger and Shallit [8] studied another class of infinite words with low complexity, namely, the Rote words, which are the infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$ [7]. Each paper gives an example of an infinite Rote word that avoids $(5/2)^{+}$-powers and shows that this is best possible among all Rote words. Ollinger and Shallit end their paper by observing that the Rote words that avoid $(5/2)^{+}$-powers appear to have a certain rigid structure reminiscent of the famous structure theorem of Restivo and Salemi [6] for the class of overlap-free words. In this note we show that this observation is indeed true and we obtain the precise structure theorem.
2. Preliminaries
For a positive integer $n$, let $\Sigma_{n}=\{0,1,\ldots,n-1\}$ and let $\Sigma_{n}^{*}$ denote the set of all finite words over $\Sigma_{n}$. By a binary word we mean a (finite or infinite) word over $\Sigma_{2}$. Let $w$ be a word and write $w=xyz$. Then the word $x$ is a prefix of $w$, the word $y$ is a factor of $w$, and the word $z$ is a suffix of $w$. A map $f:\Sigma_{m}^{*}\to\Sigma_{n}^{*}$ is a morphism if $f(xy)=f(x)f(y)$ for all $x$ and $y$.
Let $w$ have length $\ell$ and smallest period $p$. The exponent of $w$ is the quantity $k=\ell/p$ and $w$ is called a $k$-power. A $k^{+}$-power is a word with exponent $>k$. A $2$-power is a square and a $2^{+}$-power is an overlap. A word $w$ avoids $k$-powers (resp. $k^{+}$-powers) if none of its factors are $k^{\prime}$-powers for any $k^{\prime}>=k$ (resp. $k^{\prime}>k$); we also say that $w$ is $k$-power-free (resp. $k^{+}$-power-free).
Let ${\bf x}$ be an infinite word. The factor complexity of ${\bf x}$ is the function of $n$ that associates each length $n$ with the number of factors of ${\bf x}$ of length $n$. A Sturmian word is any infinite word with factor complexity $n+1$; a Rote word is any infinite word with factor complexity $2n$.
Definition 1.
Let $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ be the morphism given by
$g(0)$ $=011$
$g(1)$ $=0$
$g(2)$ $=01.$
Let $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ be the morphism given by
$f(0)$ $=0121$
$f(1)$ $=021$
$f(2)$ $=01.$
Recently the morphisms $f$ and $g$ have proved useful in several constructions [3, 5, 8]. Dvořáková et al. showed that $g(f^{\omega}(0))$ avoids $(5/2)^{+}$-powers [5], while Ollinger and Shallit showed that the factor complexity of this word is $2n$ (i.e., that it is a Rote word) [8]. The latter authors conjectured that there is a structure theorem involving $f$ and $g$ for the class of $(5/2)^{+}$-power-free Rote words.
The prototypical structure theorem of this type was obtained by Restivo and Salemi [6] for the class of binary overlap-free words. In this case, the structure is specified using the Thue–Morse morphism
$\mu(0)$ $=01$
$\mu(1)$ $=10.$
Restivo and Salemi showed the following:
Theorem 1.
Let ${\boldsymbol{w}}\in\{0,1\}^{\omega}$ be overlap-free. Then a final segment of ${\boldsymbol{w}}$ has the form $\mu({\boldsymbol{u}})$ where ${\boldsymbol{u}}$ is overlap-free.
Karhumäki and Shallit [4] later showed that the same structure theorem holds for the class of binary $(7/3)^{+}$-power-free words. In this note we establish a similar structure theorem for the class of $(5/2)^{+}$-power-free Rote words in terms of the morphisms $f$ and $g$ given above. The reader may also compare the present structure theorem with the main result of [2], which establishes a structure theorem for the class of infinite $15/4$-power-free binary rich words (rich means that every factor of length $n$ contains $n$ distinct non-empty palindromes), and which also involves a sub-family of Rote words, namely the complementary symmetric Rote words.
3. Obtaining the structure theorem
Lemma 1.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Then both of the words $0110$ and $1001$ are factors of ${\boldsymbol{w}}$.
Proof.
A backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word not containing $0110$ has length 14. Thus $0110$ (and symmetrically, $1001$) is a factor of ${\boldsymbol{w}}$. ∎
Lemma 2.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. At least 3 of the words in
$C=\{0010,0100,1011,1101\}$
are factors of ${\boldsymbol{w}}$.
Proof.
Six backtrack searches (one for each pair) show that the longest $\frac{5}{2}^{+}$ power free binary word omitting a pair of these factors has length 44. ∎
Lemma 3.
Let ${\boldsymbol{w}}$ be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. Both of $0011$ and $1100$ are factors of ${\boldsymbol{w}}$.
Proof.
Consider the set of seven binary words
$A=\{0010,0100,0101,1010,1011,1101,1100\}.$
For each word $a\in A$, a backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word containing neither of $a$ and $0011$ as a factor has length no more than 31.
It follows that if $0011$ is not a factor of ${\boldsymbol{w}}$, then ${\boldsymbol{w}}$ contains the seven words of $A$ as length 4 factors. By Lemma 1, it also contains $0110$ and $1001$ as factors. However, now ${\boldsymbol{w}}$ contains 9 factors of length 4, contradicting the fact that its factor complexity is at most $2n$.
We conclude that $0011$ (and symmetrically, $1100$) is a factor of ${\boldsymbol{w}}$. ∎
Lemma 4.
Let ${\boldsymbol{w}}$ be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. At least one of $0101$ and $1010$ is a factor of ${\boldsymbol{w}}$.
Proof.
Consider the set $D$ containing 17 binary words of length 9 given by
$D=\{00100110,01001100,10011001,00110010,01100100,11001001,$
$10010011,00110011,01100110,11001101,10011011,00110110,$
$101100 , 1$ 10110010 , 10110011 , 11001100 } .
For each word $d\in D$, a backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word containing none of $d$, $0101$, and $1010$ as a factor has length no more than 88. It follows that if neither of $0101$ and $1010$ is a factor of ${\boldsymbol{w}}$, then every word of $D$ is a factor. However, this would imply that ${\boldsymbol{w}}$ contained 17 factors of length 8, contradicting the fact that its factor complexity is at most $2n$. We conclude that at least one of $0101$ and $1010$ is a factor of ${\boldsymbol{w}}$. ∎
Theorem 2.
Let ${\boldsymbol{w}}$ be an infinite binary word with factor complexity at most $2n$, which avoids $\frac{5}{2}^{+}$ powers. Up to binary complement and/or reversal, the set of length 4 factors of ${\boldsymbol{w}}$ is
$\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Proof.
By Lemma 1, the set of length 4 factors includes $0110$ and $1001$. By Lemma 3, the set of length 4 factors includes $0011$ and $1100$. Combining Lemmas 2 and 4 with the fact that ${\boldsymbol{w}}$ has at most 8 length 4 factors, the set of length 4 factors contains exactly 3 words from $C=\{0100,0010,1011,1101\}$ and exactly one word from $\{0101,1010\}$. Since each word of $C$ maps to each of the others under complement and/or reversal, assume without loss of generality that the 3 words from $C$ are $0010$, $0100$, and $1101$. Thus ${\boldsymbol{w}}$ does not contain the factor $1011$.
A backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word not containing $1011$ or $1010$ as a factor has length 20. We conclude that ${\boldsymbol{w}}$ contains the factor $1010$, so that set of length 4 factors of ${\boldsymbol{w}}$ is
$\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
For $u\in\Sigma_{3}^{*}$, denote the Parikh vector of $u$ by $\pi(u)$. Thus,
$\pi(u)=[|u|_{0},|u|_{1},|u|_{2}].$
For $x,y\in\Sigma_{3}^{*}$, we say that $\pi(x)>\pi(y)$ if:
  1. (1)
    We have $|x|_{i}\geq|y|_{i}$ for all $i\in\{0,1,2\}$.
  2. (2)
    For at least one $i\in\{0,1,2\}$ we have $|x|_{i}>|y|_{i}.$
Lemma 5.
Suppose that $u\in\Sigma_{3}^{*}$, $w\in\Sigma_{2}^{*}$, and $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ is a non-erasing morphism. If $g(u)=w$, and $w$ avoids $\frac{5}{2}^{+}$ powers, then $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$.
Proof.
Suppose $u$ has a factor $xyxyx$ where $\pi(x)>\pi(y)$. Then $|g(x)|>|g(y)|$, and $w$ contains the $\frac{5}{2}^{+}$ power $g(x)g(y)g(x)g(y)g(x)$. ∎
Lemma 6.
Suppose that $u,v\in\Sigma_{3}^{*}$, and $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ is a non-erasing morphism. If $f(v)=u$, and $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$, then $v$ has no factor $XYXYX$ where $\pi(X)>\pi(Y)$.
Proof.
Suppose $v$ has a factor $XYXYX$ where $\pi(X)>\pi(Y)$. Let $x=f(X)$ and $y=f(Y)$. Then $\pi(x)>\pi(y)$, and $u$ contains the factor $xyxyx$. ∎
Definition 2.
Call a word $u\in\Sigma_{3}^{*}$ proper if:
  1. (1)
    Word $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$.
  2. (2)
    None of the words $00$, $11$, $22$, $20$, $10101$, $2121$, or $10210210$ is a factor of $u$.
Call a word ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ proper if all of its finite factors are proper.
Lemma 7.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Suppose that the set of length 4 factors of ${\boldsymbol{w}}$ is
$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Then a final segment of ${\boldsymbol{w}}$ has the form $g({\boldsymbol{u}})$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
Proof.
Since $111$ is not a factor of ${\boldsymbol{w}}$, any final segment of ${\boldsymbol{w}}$ beginning with 00 has the form $g({\boldsymbol{u}})$ for some ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$. We will show, replacing ${\boldsymbol{u}}$ by one of its final segments if necessary, that ${\boldsymbol{u}}$ is proper.
The fact that ${\boldsymbol{u}}$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$ follows from Lemma 5.
We conclude by showing that none of the words $00$, $11$, $22$, $20$, $10101$, $2121$, or $10210210$ is a factor of ${\boldsymbol{u}}$:
Word $00$: If $00$ is a factor of ${\boldsymbol{u}}$, then ${\boldsymbol{w}}$ contains the factor $g(00)=011011$. However, then ${\boldsymbol{w}}$ contains $1011$, which is not in $F$.
Word $11$: If $11$ is a factor of ${\boldsymbol{u}}$, then $11a$ is a factor of ${\boldsymbol{u}}$ for some $a\in\Sigma_{3}$. Then ${\boldsymbol{w}}$ contains the factor $g(11a)$, which starts with $g(1)g(1)0=000$, a $\frac{5}{2}^{+}$ power.
Word $22$: If $22$ is a factor of ${\boldsymbol{u}}$, then ${\boldsymbol{w}}$ contains the factor $g(22)=0101\notin F$.
Word $20$: If $20$ is a factor of ${\boldsymbol{u}}$, then ${\boldsymbol{w}}$ contains the factor $g(20)=01011$, which starts with $0101\notin F$.
Word $10101$: If $10101$ is not a factor of ${\boldsymbol{u}}$ more than once, replace ${\boldsymbol{u}}$ by one of its final segments not containing $10101$.
On the other hand, if $10101$ is a factor of ${\boldsymbol{u}}$ more than once, then $a10101b$ is a factor of ${\boldsymbol{u}}$ for some $a,b\in\Sigma_{3}$. Since $11$ is not a factor of ${\boldsymbol{u}}$, we can in fact specify that $a\in\Sigma_{3}-\{1\}.$ This implies that $1$ is the last letter of $g(a)$. Also, 00 is the first letter of $g(b)$. Then ${\boldsymbol{w}}$ contains the factor $g(a10101b)$, which contains $100110011001g(10101)0=$ 10011001100, a $\frac{5}{2}^{+}$ power.
Word $2121$: If $2121$ is a factor of ${\boldsymbol{u}}$, then $2121a$ is a factor of ${\boldsymbol{u}}$ for some $a\in\Sigma_{3}-\{1\}$. Thus $01$ is a prefix of $g(a)$, and ${\boldsymbol{w}}$ contains the factor $g(2121)01=01001001$, a $\frac{5}{2}^{+}$ power.
[…]
Lemma 8.
Let ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ be proper. Then a final segment of ${\boldsymbol{u}}$ has the form $f({\boldsymbol{v}})$ for some proper ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
Proof.
Assume without loss of generality, replacing ${\boldsymbol{u}}$ by a final segment if necessary, that ${\boldsymbol{u}}$ starts with the letter 00. Write ${\boldsymbol{u}}$ as a concatenation of 00-blocks, i.e., words which start with 00, and contain the letter 00 exactly once. Since ${\boldsymbol{u}}$ is proper, it does not contain a factor $00$ or $20$. It follows that its 00-blocks end with $1$. Since every occurrence of $2$ in a proper word can only be followed by a $1$, ${\boldsymbol{u}}$ cannot contain the factor $212$; otherwise it would contain the forbidden factor $2121$. We conclude that the 00-blocks of ${\boldsymbol{u}}$ are among $01$, $021$ and $0121$. Notice that the last two of these words are always followed by a 00 in ${\boldsymbol{u}}$. We thus conclude that a final segment of ${\boldsymbol{u}}$ has the form $f({\boldsymbol{v}})$ for some ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$. We will show, replacing ${\boldsymbol{v}}$ by one of its final segments if necessary, that ${\boldsymbol{v}}$ is proper.
The fact that ${\boldsymbol{v}}$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$ follows from Lemma 6.
We conclude by showing that a final segment of ${\boldsymbol{v}}$ contains none of the words $22$, $20$, $00$, $11$, $10101$, $2121$, or $10210210$ as a factor:
Word $22$: If $22$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $22$. Otherwise, $22$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $101011f(22)=$ 10101 . This is impossible, since ${\boldsymbol{w}}$ is proper.
Word $20$: If $20$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $20$. Otherwise, $20$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $10101211f(20)=$ 1010121, which starts with $10101$. This is impossible, since ${\boldsymbol{w}}$ is proper.
Word $00$: Suppose that $00$ is a factor of ${\boldsymbol{v}}$. If $000$ is a factor of ${\boldsymbol{v}}$, then $f(000)=012101210121$ is a factor of ${\boldsymbol{u}}$. However $012101210121=xyxyx$ where $x=0121$, $y=\epsilon$, and cannot be a factor of ${\boldsymbol{u}}$. It follows that $000$ is not a factor of ${\boldsymbol{v}}$.
If $00$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $00$. Otherwise, $00$ is a factor of ${\boldsymbol{v}}$ more than once, so that $100a$ is a factor of ${\boldsymbol{v}}$ for some $a\in\Sigma_{3}$. Since $f(a)$ starts with 00, this implies that ${\boldsymbol{u}}$ has a factor $f(100a)$, starting with $021012101210$. This contains the word $xyxyx$ where $x=210$, $y=1$, which is impossible, since ${\boldsymbol{u}}$ is proper.
Word $11$: If $11$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $11$. Otherwise, $11$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $102102101f(11)0=$ 10210210, which is impossible, since ${\boldsymbol{u}}$ is proper.
Word $10101$: If $10101$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $10101$. Otherwise, $10101$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $1f(10101)0=1021012102101210210=xyxyx$, where $x=10210$ and $y=12$. This is impossible, since ${\boldsymbol{u}}$ is proper.
Word $2121$: If $2121$ is not a factor of ${\boldsymbol{v}}$ more than once, then replace ${\boldsymbol{v}}$ by one of its final segments not containing $2121$. Otherwise, $2121$ is a factor of ${\boldsymbol{v}}$ more than once. Since $22$ is not a factor of $v$, $a2121b$ is a factor of ${\boldsymbol{v}}$ for some $a,b\in\Sigma_{3}$ where $a\neq 2$. Then $f(a)$ ends in $21$ and $f(b)$ begins with 00. Thus ${\boldsymbol{u}}$ contains the factor $21f(2121)0=2101021010210=xyxyx$ where $x=210$, and $y=10$, which is impossible, since ${\boldsymbol{u}}$ is proper.
Word $10210210$: If $10210210$ is a factor of ${\boldsymbol{v}}$, then ${\boldsymbol{u}}$ contains the factor $f(10210210)=0210121010210121010210121=xyxyx$, where $x=0210121$, and $y=01$, which is impossible, since ${\boldsymbol{u}}$ is proper. ∎
Theorem 3.
Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Suppose that the set of length 4 factors of ${\boldsymbol{w}}$ is
$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
Then for every positive integer $n$, a final segment of ${\boldsymbol{w}}$ has the form $g(f^{n}({\boldsymbol{u}}))$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
Proof.
This follows from Lemmas 8 and 7 by induction. ∎
Many open problems remain concerning the relationship between low factor complexity and avoidable powers. As an example, moving to a ternary alphabet, Shallit and Shur [9] showed that the word
[…]
generated by the morphism $0\to 01,1\to 2,2\to 02$, has critical exponent $2.4808627\cdots$ and factor complexity $12n+$ 1 for all $n\geq 1$. They conjecture that this exponent is minimal among all infinite ternary words with complexity $12n+$ 1. This conjecture remains open. If confirmed, it would be natural to explore the possibility of a structure theorem for this class of words.

### v3

@@
 Low complexity binary words avoiding $(5/2)^{+}$-powers
-Abstract.
+Abstract
 Rote words are infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$.
@@
 In this note we give a structure theorem for the Rote words that avoid $(5/2)^{+}$-powers, confirming a conjecture of Ollinger and Shallit.
-The work of James Currie is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number DDG-2024-00005].
-The work of Narad Rampersad is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2019-04111].
-1. Introduction
+keywords:
+Rote word, factor complexity, $(5/2)^{+}$-power, structure theorem
+1 Introduction
 Two central concepts in combinatorics on words are power avoidance and factor complexity.
-Recently, Shallit and Shur [9] initiated the systematic investigation of the interplay between these two concepts.
+Recently, Shallit and Shur (2019) initiated the systematic investigation of the interplay between these two concepts.
 They examined two dual problems: 1) Given a particular power avoidance constraint, determine the range of possible factor complexities among all infinite words avoiding the specified power; and, 2) Given a class of words with specified factor complexities, determine the powers that are avoided by at least one word in this class.
-A well-known classical result provides a solution for the latter problem for the class of Sturmian words, i.e., the class of infinite words that contain $n+1$ factors of length $n$ for every $n\geq 1$: the Fibonacci word avoids $(5+\sqrt{5})/2$-powers, and this is best possible among all Sturmian words [1].
-The Sturmian words are the aperiodic infinite words with the least possible factor complexity function; Shallit and Shur [9] and Ollinger and Shallit [8] studied another class of infinite words with low complexity, namely, the Rote words, which are the infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$ [7].
+A well-known classical result provides a solution for the latter problem for the class of Sturmian words, i.e., the class of infinite words that contain $n+1$ factors of length $n$ for every $n\geq 1$: the Fibonacci word avoids $(5+\sqrt{5})/2$-powers, and this is best possible among all Sturmian words (Carpi and de Luca (2000)).
+The Sturmian words are the aperiodic infinite words with the least possible factor complexity function; Shallit and Shur (2019) and Ollinger and Shallit (2024) studied another class of infinite words with low complexity, namely, the Rote words, which are the infinite words that contain $2n$ factors of length $n$ for every $n\geq 1$ (Rote (1994)).
 Each paper gives an example of an infinite Rote word that avoids $(5/2)^{+}$-powers and shows that this is best possible among all Rote words.
-Ollinger and Shallit end their paper by observing that the Rote words that avoid $(5/2)^{+}$-powers appear to have a certain rigid structure reminiscent of the famous structure theorem of Restivo and Salemi [6] for the class of overlap-free words.
-In this note we show that this observation is indeed true and we obtain the precise structure theorem.
-2. Preliminaries
+Ollinger and Shallit end their paper by observing that the Rote words that avoid $(5/2)^{+}$-powers appear to have a certain rigid structure reminiscent of the famous structure theorem of Restivo and Salemi (1985) for the class of overlap-free words.
+In this note we obtain the precise structure theorem.
+To state our first structure theorem, we introduce proper words and antiproper words, which are ternary words.
+Definition 1.
+For $u\in\Sigma_{3}^{*}$, denote the Parikh vector of $u$ by $\pi(u)$so that $\pi(u)=[|u|_{0},|u|_{1},|u|_{2}].$ For $x,y\in\Sigma_{3}^{*}$, we say that $\pi(x)>\pi(y)$ if:
+  1. 1.
+    We have $|x|_{i}\geq|y|_{i}$ for all $i\in\Sigma_{3}$.
+  2. 2.
+    For at least one $i\in\Sigma_{3}$ we have $|x|_{i}>|y|_{i}.$
+Call a word $u\in\Sigma_{3}^{*}$ proper if:
+  1. 1.
+    Word $u$ has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$.
+  2. 2.
+    None of the words $00$, $11$, $22$, 2020, $10101$, $2121$, or $10210210$ is a factor of $u$.
+Call a word ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ proper if all of its finite factors are proper.
+Call a word $u\in\Sigma_{3}^{*}$ antiproper if its reverse $u^{R}$ is proper. Call a word ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ antiproper if all of its finite factors are antiproper.
+Proper words obey a structure theorem similar to that of Restivo and Salemi, as do antiproper words.
+We introduce a morphism $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ and its reverse $h:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$, given by:
+$f(0)$ $=0121$ $h(0)$ $=1210$
+$f(1)$ $=021$ $h(1)$ $=120$
+$f(2)$ $=01$ $h(2)$ $=10.$
+Theorem 1.
+(First Structure Theorem)
+  1. 1.
+    Let ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ be proper. Then a final segment of $\boldsymbol{u}$} has the form $f({\boldsymbol{v}})$ for some proper ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
+  2. 2.
+    Let ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ be antiproper.
+Then a final segment of $\boldsymbol{u}$} has the form $h({\boldsymbol{v}})$ for some antiproper ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
+For our second structure theorem we consider the length-4 factors of a Rote word.
+Definition 2.
+Let $F$ be the set
+$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
+Let $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ be the morphism given by
+$g(0)$ $=011$
+$g(1)$ $=0$
+$g(2)$ $=01.$
+We denote the complement of a binary word $w$ by $\overline{w}$; thus $\overline{1101}=0010$.
+We extend this notation to binary languages in the usual way.
+We denote the reversal of a word $w$ by $w^{R}$; thus $1101^{R}=1011$.
+We extend this notation to languages in the usual way.
+We can now characterize the structure of Rote words that avoid $(5/2)^{+}$-powers.
+Theorem 2.
+(Second Structure Theorem) Let $\boldsymbol{w}$} be a Rote word that avoids $(5/2)^{+}$-powers. There are four cases:
+  1. 1.
+    The set of length-4 factors of $\boldsymbol{w}$} is $F$.
+For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $g(f^{n}({\boldsymbol{u}}))$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
+  2. 2.
+    The set of length-4 factors of $\boldsymbol{w}$} is $\bar{F}$.
+For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $\overline{g(f^{n}({\boldsymbol{u}}))}$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
+  3. 3.
+    The set of length-4 factors of $\boldsymbol{w}$} is $F^{R}$.
+For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $g(h^{n}({\boldsymbol{u}}))$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
+  4. 4.
+    The set of length-4 factors of $\boldsymbol{w}$} is $\overline{F^{R}}$.
+For every positive integer $n$, a final segment of $\boldsymbol{w}$} has the form $\overline{g(h^{n}({\boldsymbol{u}}))}$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
+2 Preliminaries
 For a positive integer $n$, let $\Sigma_{n}=\{0,1,\ldots,n-1\}$ and let $\Sigma_{n}^{*}$ denote the set of all finite words over $\Sigma_{n}$.
@@
 The exponent of $w$ is the quantity $k=\ell/p$ and $w$ is called a $k$-power.
-A $k^{+}$-power is a word with exponent $>k$.
+A $k^{+}$-power is a word with exponent >k>k.
 A $2$-power is a square and a $2^{+}$-power is an overlap.
@@
 A Sturmian word is any infinite word with factor complexity $n+1$; a Rote word is any infinite word with factor complexity $2n$.
-Definition 1.
-Let $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ be the morphism given by
-$g(0)$ $=011$
-$g(1)$ $=0$
-$g(2)$ $=01.$
-Let $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ be the morphism given by
-$f(0)$ $=0121$
-$f(1)$ $=021$
-$f(2)$ $=01.$
-Recently the morphisms $f$ and $g$ have proved useful in several constructions [3, 5, 8].
+Recently the morphisms $f$ and $g$ have proved useful in several constructions (Currie et al.
+(2023); Dvořáková et al.
+(2024); Ollinger and Shallit (2024)).
 Dvořáková et al.
-showed that $g(f^{\omega}(0))$ avoids $(5/2)^{+}$-powers [5], while Ollinger and Shallit showed that the factor complexity of this word is $2n$ (i.e., that it is a Rote word) [8].
+(2024) showed that $g(f^{\omega}(0))$ avoids $(5/2)^{+}$-powers, while Ollinger and Shallit (2024) showed that the factor complexity of this word is $2n$ (i.e., that it is a Rote word).
 The latter authors conjectured that there is a structure theorem involving $f$ and $g$ for the class of $(5/2)^{+}$-power-free Rote words.
-The prototypical structure theorem of this type was obtained by Restivo and Salemi [6] for the class of binary overlap-free words.
-In this case, the structure is specified using the Thue–Morse morphism
+A prototypical structure theorem of this type was obtained by Restivo and Salemi (1985) for the class of binary overlap-free words.
+(See also Fife (1980); Shur (1996); Shallit (2011).) In this case, the structure is specified using the Thue–Morse morphism
 $\mu(0)$ $=01$
@@
 Theorem 3.
-Let ${\boldsymbol{w}}\in\{0,1\}^{\omega}$ be overlap-free. Then a final segment of ${\boldsymbol{w}}$ has the form $\mu({\boldsymbol{u}})$ where ${\boldsymbol{u}}$ is overlap-free.
-Karhumäki and Shallit [4] later showed that the same structure theorem holds for the class of binary $(7/3)^{+}$-power-free words.
+Let ${\boldsymbol{w}}\in\Sigma_{2}^{\omega}$ be overlap-free. Then a final segment of $\boldsymbol{w}$} has the form $\mu({\boldsymbol{u}})$ where $\boldsymbol{u}$} is overlap-free.
+Karhumäki and Shallit (2004) later showed that the same structure theorem holds for the class of binary $(7/3)^{+}$-power-free words.
 In this note we establish a similar structure theorem for the class of $(5/2)^{+}$-power-free Rote words in terms of the morphisms $f$ and $g$ given above.
-The reader may also compare the present structure theorem with the main result of [2], which establishes a structure theorem for the class of infinite $15/4$-power-free binary rich words (rich means that every factor of length $n$ contains $n$ distinct non-empty palindromes), and which also involves a sub-family of Rote words, namely the complementary symmetric Rote words.
-3. Obtaining the structure theorem
+The reader may also compare the present structure theorem with the main result of Currie et al.
+(2020–2021), which establishes a structure theorem for the class of infinite $14/5$-power-free binary rich words (rich means that every factor of length $n$ contains $n$ distinct non-empty palindromes), and which also involves a sub-family of Rote words, namely the complementary symmetric Rote words.
+3 Obtaining the structure theorem
 Lemma 1.
@@
 $D=\{00100110,01001100,10011001,00110010,01100100,11001001,$
-$10010011,00110011,01100110,11001101,10011011,00110110,$
-$101100 , 1$ 10110010 , 10110011 , 11001100 } .
+10010011,00110011,01100110,11001101,10011011,00110110,10010011,00110011,01100110,11001101,10011011,00110110,
+01101100,11011001,10110010,10110011,11001100}.01101100,11011001,10110010,10110011,11001100\}.
 For each word $d\in D$, a backtrack search shows that the longest $\frac{5}{2}^{+}$ power free binary word containing none of $d$, $0101$, and $1010$ as a factor has length no more than 88.
@@
 $\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
-For $u\in\Sigma_{3}^{*}$, denote the Parikh vector of $u$ by $\pi(u)$. Thus,
-$\pi(u)=[|u|_{0},|u|_{1},|u|_{2}].$
-For $x,y\in\Sigma_{3}^{*}$, we say that $\pi(x)>\pi(y)$ if:
-  1. (1)
-    We have $|x|_{i}\geq|y|_{i}$ for all $i\in\{0,1,2\}$.
-  2. (2)
-    For at least one $i\in\{0,1,2\}$ we have $|x|_{i}>|y|_{i}.$
 Lemma 5.
 Suppose that $u\in\Sigma_{3}^{*}$, $w\in\Sigma_{2}^{*}$, and $g:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ is a non-erasing morphism.
-If $g(u)=w$, and $w$ avoids $\frac{5}{2}^{+}$ powers, then $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$.
+If $g(u)=w$, and $w$ avoids $\frac{5}{2}^{+}$ powers, then $u$ has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$.
 Proof.
-Suppose $u$ has a factor $xyxyx$ where $\pi(x)>\pi(y)$. Then $|g(x)|>|g(y)|$, and $w$ contains the $\frac{5}{2}^{+}$ power $g(x)g(y)g(x)g(y)g(x)$. ∎
+Suppose $u$ has a factor xyxyxxyxyx where $\pi(x)>\pi(y)$. Then $|g(x)|>|g(y)|$, and $w$ contains the $\frac{5}{2}^{+}$ power $g(x)g(y)g(x)g(y)g(x)$. ∎
 Lemma 6.
 Suppose that $u,v\in\Sigma_{3}^{*}$, and $f:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ is a non-erasing morphism.
-If $f(v)=u$, and $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$, then $v$ has no factor $XYXYX$ where $\pi(X)>\pi(Y)$.
+If $f(v)=u$, and $u$ has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$, then $v$ has no factor $XYXYX$ where $\pi(X)>\pi(Y)$.
 Proof.
-Suppose $v$ has a factor $XYXYX$ where $\pi(X)>\pi(Y)$. Let $x=f(X)$ and $y=f(Y)$. Then $\pi(x)>\pi(y)$, and $u$ contains the factor $xyxyx$. ∎
-Definition 2.
-Call a word $u\in\Sigma_{3}^{*}$ proper if:
-  1. (1)
-    Word $u$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$.
-  2. (2)
-    None of the words $00$, $11$, $22$, $20$, $10101$, $2121$, or $10210210$ is a factor of $u$.
-Call a word ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ proper if all of its finite factors are proper.
+Suppose $v$ has a factor $XYXYX$ where $\pi(X)>\pi(Y)$. Let $x=f(X)$ and $y=f(Y)$. Then $\pi(x)>\pi(y)$, and $u$ contains the factor xyxyxxyxyx. ∎
 Lemma 7.
@@
 Proof.
-Since $111$ is not a factor of ${\boldsymbol{w}}$, any final segment of ${\boldsymbol{w}}$ beginning with 00 has the form $g({\boldsymbol{u}})$ for some ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
+Since $111$ is not a factor of $\boldsymbol{w}$}, any final segment of $\boldsymbol{w}$} beginning with 0 has the form $g({\boldsymbol{u}})$ for some ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
 We will show, replacing $\boldsymbol{u}$} by one of its final segments if necessary, that $\boldsymbol{u}$} is proper.
-The fact that ${\boldsymbol{u}}$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$ follows from Lemma 5.
-We conclude by showing that none of the words $00$, $11$, $22$, $20$, $10101$, $2121$, or $10210210$ is a factor of ${\boldsymbol{u}}$:
+The fact that $\boldsymbol{u}$} has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$ follows from Lemma 5.
+We conclude by showing that none of the words $00$, $11$, $22$, 2020, $10101$, $2121$, or $10210210$ is a factor of $\boldsymbol{u}$}:
 Word $00$: If $00$ is a factor of $\boldsymbol{u}$}, then $\boldsymbol{w}$} contains the factor $g(00)=011011$. However, then $\boldsymbol{w}$} contains $1011$, which is not in $F$.
@@
 Since $11$ is not a factor of $\boldsymbol{u}$}, we can in fact specify that $a\in\Sigma_{3}-\{1\}.$ This implies that $1$ is the last letter of $g(a)$.
-Also, 00 is the first letter of $g(b)$.
-Then ${\boldsymbol{w}}$ contains the factor $g(a10101b)$, which contains $100110011001g(10101)0=$ 10011001100, a $\frac{5}{2}^{+}$ power.
+Also, 0 is the first letter of $g(b)$.
+Then $\boldsymbol{w}$} contains the factor $g(a10101b)$, which contains $1g(10101)0=10011001100$, a $\frac{5}{2}^{+}$ power.
 Word $2121$: If $2121$ is a factor of $\boldsymbol{u}$}, then $2121a$ is a factor of $\boldsymbol{u}$} for some $a\in\Sigma_{3}-\{1\}$.
 Thus $01$ is a prefix of $g(a)$, and $\boldsymbol{w}$} contains the factor $g(2121)01=01001001$, a $\frac{5}{2}^{+}$ power.
-[…]
-Lemma 8.
-Let ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$ be proper. Then a final segment of ${\boldsymbol{u}}$ has the form $f({\boldsymbol{v}})$ for some proper ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
-Proof.
-Assume without loss of generality, replacing ${\boldsymbol{u}}$ by a final segment if necessary, that ${\boldsymbol{u}}$ starts with the letter 00.
-Write ${\boldsymbol{u}}$ as a concatenation of 00-blocks, i.e., words which start with 00, and contain the letter 00 exactly once.
+Word $10210210$: If $10210210$ is a factor of $\boldsymbol{u}$}, then $\boldsymbol{w}$} contains the factor
+$g(10210210)=0011010011010011,$
+a $\frac{5}{2}^{+}$ power. ∎
+We can now prove Theorem 1:
+of Theorem 1.
+Assume without loss of generality, replacing $\boldsymbol{u}$} by a final segment if necessary, that $\boldsymbol{u}$} starts with the letter 0.
+Write $\boldsymbol{u}$} as a concatenation of 0-blocks, i.e., words which start with 0, and contain the letter 0 exactly once.
 Since $\boldsymbol{u}$} is proper, it does not contain a factor $00$ or $20$.
-It follows that its 00-blocks end with $1$.
+It follows that its 0-blocks end with $1$.
 Since every occurrence of $2$ in a proper word can only be followed by a $1$, $\boldsymbol{u}$} cannot contain the factor $212$; otherwise it would contain the forbidden factor $2121$.
-We conclude that the 00-blocks of ${\boldsymbol{u}}$ are among $01$, $021$ and $0121$.
-Notice that the last two of these words are always followed by a 00 in ${\boldsymbol{u}}$.
+We conclude that the 0-blocks of $\boldsymbol{u}$} are among $01$, $021$ and $0121$.
+Notice that the last two of these words are always followed by a 0 in $\boldsymbol{u}$}.
 We thus conclude that a final segment of $\boldsymbol{u}$} has the form $f({\boldsymbol{v}})$ for some ${\boldsymbol{v}}\in\Sigma_{3}^{\omega}$.
 We will show, replacing $\boldsymbol{v}$} by one of its final segments if necessary, that $\boldsymbol{v}$} is proper.
-The fact that ${\boldsymbol{v}}$ has no factor $xyxyx$ where $\pi(x)>\pi(y)$ follows from Lemma 6.
-We conclude by showing that a final segment of ${\boldsymbol{v}}$ contains none of the words $22$, $20$, $00$, $11$, $10101$, $2121$, or $10210210$ as a factor:
+The fact that $\boldsymbol{v}$} has no factor xyxyxxyxyx where $\pi(x)>\pi(y)$ follows from Lemma 6.
+We conclude by showing that a final segment of $\boldsymbol{v}$} contains none of the words $22$, 2020, $00$, $11$, $10101$, $2121$, or $10210210$ as a factor:
 Word $22$: If $22$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $22$.
-Otherwise, $22$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $101011f(22)=$ 10101 .
-This is impossible, since ${\boldsymbol{w}}$ is proper.
+Otherwise, $22$ is a factor of $\boldsymbol{v}$} more than once, so that $\boldsymbol{u}$} contains a factor $1f(22)=10101.$ This is impossible, since $\boldsymbol{u}$} is proper.
 Word $20$: If $20$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $20$.
-Otherwise, $20$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $10101211f(20)=$ 1010121, which starts with $10101$.
-This is impossible, since ${\boldsymbol{w}}$ is proper.
+Otherwise, $20$ is a factor of $\boldsymbol{v}$} more than once, so that $\boldsymbol{u}$} contains a factor $1f(20)=1010121$, which starts with $10101$.
+This is impossible, since $\boldsymbol{u}$} is proper.
 Word $00$: Suppose that $00$ is a factor of $\boldsymbol{v}$}.
@@
 Otherwise, $00$ is a factor of $\boldsymbol{v}$} more than once, so that $100a$ is a factor of $\boldsymbol{v}$} for some $a\in\Sigma_{3}$.
-Since $f(a)$ starts with 00, this implies that ${\boldsymbol{u}}$ has a factor $f(100a)$, starting with $021012101210$.
-This contains the word $xyxyx$ where $x=210$, $y=1$, which is impossible, since ${\boldsymbol{u}}$ is proper.
+Since $f(a)$ starts with 0, this implies that $\boldsymbol{u}$} has a factor $f(100a)$, starting with $021012101210$.
+This contains the word xyxyxxyxyx where $x=210$, $y=1$, which is impossible, since $\boldsymbol{u}$} is proper.
 Word $11$: If $11$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $11$.
-Otherwise, $11$ is a factor of ${\boldsymbol{v}}$ more than once, so that ${\boldsymbol{u}}$ contains a factor $102102101f(11)0=$ 10210210, which is impossible, since ${\boldsymbol{u}}$ is proper.
+Otherwise, $11$ is a factor of $\boldsymbol{v}$} more than once, so that $\boldsymbol{u}$} contains a factor $1f(11)0=10210210$, which is impossible, since $\boldsymbol{u}$} is proper.
 Word $10101$: If $10101$ is not a factor of $\boldsymbol{v}$} more than once, then replace $\boldsymbol{v}$} by one of its final segments not containing $10101$.
@@
 Since $22$ is not a factor of $v$, $a2121b$ is a factor of $\boldsymbol{v}$} for some $a,b\in\Sigma_{3}$ where $a\neq 2$.
-Then $f(a)$ ends in $21$ and $f(b)$ begins with 00.
+Then $f(a)$ ends in $21$ and $f(b)$ begins with 0.
 Thus $\boldsymbol{u}$} contains the factor $21f(2121)0=2101021010210=xyxyx$ where $x=210$, and $y=10$, which is impossible, since $\boldsymbol{u}$} is proper.
-Word $10210210$: If $10210210$ is a factor of ${\boldsymbol{v}}$, then ${\boldsymbol{u}}$ contains the factor $f(10210210)=0210121010210121010210121=xyxyx$, where $x=0210121$, and $y=01$, which is impossible, since ${\boldsymbol{u}}$ is proper.
-∎
-Theorem 3.
-Let ${\boldsymbol{w}}$ be an infinite binary word which avoids $\frac{5}{2}^{+}$ powers. Suppose that the set of length 4 factors of ${\boldsymbol{w}}$ is
-$F=\{0110,1001,0011,1100,0010,0100,1101,1010\}.$
-Then for every positive integer $n$, a final segment of ${\boldsymbol{w}}$ has the form $g(f^{n}({\boldsymbol{u}}))$ for some proper ${\boldsymbol{u}}\in\Sigma_{3}^{\omega}$.
-Proof.
-This follows from Lemmas 8 and 7 by induction. ∎
-Many open problems remain concerning the relationship between low factor complexity and avoidable powers. As an example, moving to a ternary alphabet, Shallit and Shur [9] showed that the word
-[…]
-generated by the morphism $0\to 01,1\to 2,2\to 02$, has critical exponent $2.4808627\cdots$ and factor complexity $12n+$ 1 for all $n\geq 1$.
-They conjecture that this exponent is minimal among all infinite ternary words with complexity $12n+$ 1.
-This conjecture remains open.
-If confirmed, it would be natural to explore the possibility of a structure theorem for this class of words.
+Word $10210210$: If $10210210$ is a factor of $\boldsymbol{v}$}, then $\boldsymbol{u}$} contains the factor
+f(10210210)=0210121010210121010210121=xyxyx,f(10210210)=0210121010210121010210121=xyxyx,
+where $x=0210121$, and $y=01$, which is impossible, since $\boldsymbol{u}$} is proper.
+The proof for antiproper words is the same, mutatis mutandi. ∎
+We can now prove Theorem 2:
+of Theorem 2.
+The first case follows from Theorem 1 and Lemma 7 by induction. The other cases follow, mutatis mutandi. ∎
+As an example of the third case of the theorem, Shallit and Shur (2019) consider a word $\tau({\bf G})$, where $\tau:\Sigma_{3}^{*}\rightarrow\Sigma_{2}^{*}$ is the morphism given by
+$\tau(0)$ $=0$
+$\tau(1)$ $=01$
+$\tau(2)$ $=011$
+and $\bf G$} is the fixed point of $\theta$, where $\theta:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ is the morphism given by
+$\theta(0)$ $=01$
+$\theta(1)$ $=2$
+$\theta(2)$ $=02.$
+Letting $\sigma$ be the permutation $0\to 1\to 2\to 0$, one checks that $\tau=g\sigma$ and $\theta^{2}=\sigma^{-1}h\sigma$. It follows that
+$\tau({\bf G})$ $=\tau(\theta^{2n}(\theta^{\omega}(0)))$
+$=g\sigma((\sigma^{-1}h\sigma)^{n-1}(\sigma^{-1}h\sigma(\theta^{\omega}(0)))$
+$=gh^{n-1}({\boldsymbol{u}})),$
+where ${\boldsymbol{u}}=h(\sigma(\theta^{\omega}(0)))$, which is antiproper.
+Many open problems remain concerning the relationship between low factor complexity and avoidable powers. Moving to a ternary alphabet, Shallit and Shur (2019) showed that the word
+${\bf G}=0120201020120102012\cdots$
+has critical exponent $2.4808627\cdots$ and factor complexity $2n+1$ for all $n\geq 1$.
+They conjectured that this exponent is minimal among all infinite ternary words with complexity $2n+1$.
+This conjecture was recently confirmed by Currie (2025).
+It would thus be natural to explore the possibility of a structure theorem for this class of words.
+4 Appendix: Python code and output
+The backtrack searches mentioned in the paper run quickly in Python. Here is our code and its output:
+def fhpf(w): #Word w is 5/2+ power suffix free
+    p=1
+    while (5*p<2*len(w)):
+        if (w[(-(p+1)//2)-p:]==w[(-(p+1)//2)-2*p:-p]):
+            return(False)
+        p=p+1
+    return(True)
+def good(w): # Word w has no suffix which is a 5/2+ power, or is in the
+             # set Factors.
+    for f in Factors:
+        k=len(f)
+        if ((len(w)>= k)and(w[-k:]==f)):
+            return(False)
+    return(fhpf(w))
+def search(target): # This returns the lexicographically least word not
+                    # containing a 5/2+ power or a word in the set Factors
+    w=’’
+    Max=0
+    while (len(w)<=target):
+        if (good(w)):
+            Max=max(Max,len(w))
+            if (len(w)==target):
+                return(w)
+            w+=’0’
+        else:
+            while((len(w)>0) and (w[-1]==’1’)):
+                w=w[:-1]
+            if(len(w)==0):
+                print(’Longest 5/2+-power-free word with no factor in ’,Factors,’
+                    has length ’,Max)
+                return()
+            c=chr(ord(w[-1])+1)
+            w=w[:-1]
+            w+=c
+    return()
+# Lemma 1
+print(’=============’)
+print(’Computations for Lemma 1’)
+print(’ ’)
+Factors=[’0110’]
+search(200)
+# Lemma 2
+print(’ ’)
+print(’=============’)
+print(’Computations for Lemma 2’)
+print(’ ’)
+C=[’0010’,’0100’,’1011’,’1101’]
+for i in range(4):
+    for j in range(i,4):
+        Factors=[C[i]]
+        Factors.append(C[j])
+        search(200)
+# Lemma 3
+print(’ ’)
+print(’=============’)
+print(’Computations for Lemma 3’)
+print(’ ’)
+A=[’0010’,’0100’,’0101’,’1010’,’1011’,’1101’,’1100’]
+for j in A:
+    Factors=[’0011’]
+    Factors.append(j)
+    search(200)
+# Lemma 4
+print(’ ’)
+print(’=============’)
+print(’Computations for Lemma 4’)
+print(’ ’)
+D=[’00100110’,’01001100’,’10011001’,’00110010’,
+ ’01100100’,’11001001’,’10010011’,’$0011$’,
+ ’01100110’,’11001101’,’10011011’,’00110110’,
+ ’01101100’,’11011001’,’10110010’,’10110011’,’$1100$’]
+for j in D:
+    Factors=[’0101’,’$10$’]
+    Factors.append(j)
+    search(200)
+# Theorem 4
+print(’ ’)
+print(’=============’)
+print(’Computations for Theorem 4’)
+print(’ ’)
+Factors=[’1011’,’$10$’]
+search(200)
+Computations for Lemma 1
+Longest 5/2+-power-free word with no factor in  [’0110’]  has length  14
+Computations for Lemma 2
+Longest 5/2+-power-free word with no factor in  [’0010’, ’0100’]  has length  44
+Longest 5/2+-power-free word with no factor in  [’0010’, ’1011’]  has length  28
+Longest 5/2+-power-free word with no factor in  [’0010’, ’1101’]  has length  13
+Longest 5/2+-power-free word with no factor in  [’0100’, ’1011’]  has length  13
+Longest 5/2+-power-free word with no factor in  [’0100’, ’1101’]  has length  28
+Longest 5/2+-power-free word with no factor in  [’1011’, ’1101’]  has length  44
+Computations for Lemma 3
+Longest 5/2+-power-free word with no factor in  [’0011’, ’0010’]  has length  15
+Longest 5/2+-power-free word with no factor in  [’0011’, ’0100’]  has length  31
+Longest 5/2+-power-free word with no factor in  [’0011’, ’0101’]  has length  12
+Longest 5/2+-power-free word with no factor in  [’0011’, ’1010’]  has length  18
+Longest 5/2+-power-free word with no factor in  [’0011’, ’1011’]  has length  15
+Longest 5/2+-power-free word with no factor in  [’0011’, ’1101’]  has length  31
+Longest 5/2+-power-free word with no factor in  [’0011’, ’1100’]  has length  30
+Computations for Lemma 4
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00100110’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01001100’]  has length  50
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10011001’]  has length  33
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00110010’]  has length  50
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01100100’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11001001’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10010011’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00110011’]  has length  52
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01100110’]  has length  33
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11001101’]  has length  50
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10011011’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’00110110’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’01101100’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11011001’]  has length  24
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10110010’]  has length  88
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’10110011’]  has length  50
+Longest 5/2+-power-free word with no factor in  [’0101’, ’1010’, ’11001100’]  has length  52
+Computations for Theorem 4
+Longest 5/2+-power-free word with no factor in  [’1011’, ’1010’]  has length  20